in how many ways can six people in line get on a bus if three person insist to follow each other?

Answer by fcabanski(1391)

Since the three people want to follow each other treat that group of three as one person. But first determine how many combinations those three people can take.


If they are called A, B and C the first person could be any of the three 3, while the second person could be any of the remaining 2, and the third person is the remaining one. So amongst those three people there are 3*2*1 = 6 possible orders.


With that group taken as one unit (call ABC G, the group), there are 4 choices for the first spot (G, D, E, F) 4, and then 3 choices for the second spot, 2 for the third and 1 for the final spot. So that's 4*3*2*1 = 24 possibilities.


The total is 6 * 24 = 144 possible combinations.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com